A chemist needs to produce a 60% acid mixture. She has 25% strength solution and 80% strength solution on hand.
Rounding to the nearest thousandth (3 decimal places; nearest mL), how many liters of 80% solution does she add to create 1 liter of the 60% strength mixture?

Found 2 solutions by stanbon, checkley79:
Answer by stanbon(75887)

A chemist needs to produce a 60% acid mixture. She has 25% strength solution and 80% strength solution on hand.
Rounding to the nearest thousandth (3 decimal places; nearest mL), how many liters of 80% solution does she add to create 1 liter of the 60% strength mixture?
--------
Equation:
acid + acid = acid
0.25x + 0.80(1-x) = 0.60*1
-------
25x + 80 - 80x = 60
-55x = -20
x = 4/11 liter (amt. of 25% solution needed)
----
1-4/11 = 7/11 liter (amt of 80% solution needed)
==============
Cheers,
Stan H.
==============

Answer by checkley79(3341)

.80x+.25(1-x)=.60*1
.80x+.25-.25x=.60
.55x=.60-.25
.55x=.35
x=.35/.55
x=.636 liters of 80% solution is used.
1-.636=.364 liters of 25% solution is used.
Proof:
.80*.636+.25*.364=.60
.51+.09=.60
.60=.60